Using relation for accessible state,(omega)=B(v-b)^Nf(E) here f(E) is function of energy only. Obtain equation of state p(v-b)=NkT
1
Expert's answer
2020-04-15T10:32:48-0400

As per the given question,

Average potential energy of the molecules

"u=\dfrac{-a}{NV}"

Free volume per molecule, "V_f=\dfrac{(V-b)}{N}"

The molecular partition function,

"z=(\dfrac{2\pi mkT}{h^2})^{\frac{3}{2}}\int_o^{v_f}e^{\frac{-u}{kT}}dv"

"\Rightarrow z=(\dfrac{2\pi mkT}{h^2})^{\frac{3}{2}}(v_f)e^{\frac{-u}{kT}}"

Now, as per the definition, substituting the values,

"\ln z=\dfrac{3}{2}\ln(\dfrac{2\pi mkT}{h^2})^{\frac{3}{2}}+\ln (V-b)+\dfrac{a}{VNkT}"

Now, free energy equation can be written as,

"F=-NkT(\dfrac{3}{2}\ln(\dfrac{2\pi mkT}{h^2})^{\frac{3}{2}}+\ln (V-b))-\dfrac{a}{V}"

The pressure of such gas containing N molecule in volume V,

"P=NkT(\dfrac{\partial \ln z}{\partial V})_T=\dfrac{NkT}{V-b}-\dfrac{a}{V^2}"

if a=0, then

"P(V-b)=NkT"